40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Mar 21 10:31:28 2012, steve, Update, SAFETY, AC power to ETMY turned off 
    Reply  Wed Mar 21 11:32:47 2012, steve, Update, SAFETY, AC power to ETMX turned off 
       Reply  Wed Mar 21 19:12:27 2012, steve, Update, Computers, AC power back on both ends 
          Reply  Thu Mar 22 16:45:06 2012, kiwamu, Update, CDS, c1scx and c1scy not properly running 
             Reply  Thu Mar 22 17:41:15 2012, suresh, Update, CDS, c1scx and c1scy not properly running 
                Reply  Thu Mar 22 23:43:56 2012, Koji, Update, CDS, c1scx and c1scy not properly running 
Message ID: 6432     Entry time: Wed Mar 21 10:31:28 2012     Reply to this: 6433
Author: steve 
Type: Update 
Category: SAFETY 
Subject: AC power to ETMY turned off 

ETMY sus damping was disabled. Green locking laser and associated electronics turned off. Computers and power supplies turned off at rack 1Y4

The electricians picking up ac power from 1Y4 manual disconnect box and installing conduit line to ISCT-ETMY east end optical table.

There will be no more daisy chaining this way. 

ELOG V3.1.3-